OpenStudy (cookiimonster627):

Expand the binomials using the Binomial theorem. (2x+3y)^4

OpenStudy (campbell_st):

binomial theorem says \[(a + b)^n = \left(\begin{matrix}n \\ r\end{matrix}\right) \sum_{n = 0}^{k}(a)^{n - k}(b)^k\]

OpenStudy (cookiimonster627):

how do I figure it out

OpenStudy (campbell_st):

ok... so a = 2x and b = 3y and n = 4 so rather than using combinations to get the initial coefficient you can just use Pascal's Triangle so its 1, 4, 6, 4, 1 so the then terms are \[1 \times (2x)^4(3y)^0 + 4 \times (2x)^4(3y)^1 + 6 \times (2x)^3 (3y)^2+...\] so as the power of (2x) decreases the power of (3y) increases... there are 5 terms... so when you have all 5 just multiply the coefficients \[16x^4 + 4 \times 8x^3\times2y+ ..... \]

OpenStudy (campbell_st):

hope it makes sense

OpenStudy (cookiimonster627):

yea thank you

OpenStudy (campbell_st):

opps I think I got the powers and sum from and to.. round the wrong way

OpenStudy (cookiimonster627):

what do you mean?